OpenStudy (anonymous):

Driving on the highway at a constant speed, it takes you 24 minutes to travel 20 miles. How long will it take you to complete a 70-mile trip?

OpenStudy (anonymous):

\[\frac{24\min}{20miles}=\frac{12\min}{10miles}=1.2\frac{\min}{mile}\]

OpenStudy (anonymous):

Now \[1.2\frac{\min}{mile} (70mile) = 84\min\]